OpenStudy (theedwardsfamily):

Medal will be rewarded!!!!!!!!!!!!! :) Evaluate. 6[5(3-9)-1]+2 ______________ 7(8-2)+4 @Here_to_Help15 @Godlovesme @dtan5457 @TheSmartOne @sleepyjess @pooja195

TheSmartOne (thesmartone):

PEMDAS.

OpenStudy (godlovesme):

first subtract 3-9

Nnesha (nnesha):

PEMDAS aunti

OpenStudy (toriii):

-4

OpenStudy (theedwardsfamily):

nice joke nnesh. @Godlovesme for the top got -184

OpenStudy (godlovesme):

hold on @TheEdwardsFamily

OpenStudy (theedwardsfamily):

kk

OpenStudy (godlovesme):

i got -22/46 do u have options tho?

OpenStudy (theedwardsfamily):

im still working it out

OpenStudy (godlovesme):

me too

OpenStudy (theedwardsfamily):

@texaschic101

OpenStudy (godlovesme):

@Directrix please help

Directrix (directrix):

[ 6 * [ 5 * ( 3 - 9 ) -1 ]+ 2 ] / [ 7 * ( 8 - 2 ) + 4] = ---------------------------------------- Numerator: [ 6 * [ 5 * ( 3 - 9 ) -1 ]+ 2 ] Denominator: [ 7 * ( 8 - 2 ) + 4]

Directrix (directrix):

Work from the inside out: Numerator: [ 6 * [ 5 * ( 3 - 9 ) -1 ]+ 2 ] [ 6 * [ 5 * ( 3 - 9 ) -1 ] + 2 ] = [ 6 * [ 5 * ( -6 ) -1 ] + 2 ] = [ 6 * [ ( - 30) -1 ] + 2 ] = [ 6 * [ -31 ] + 2 ] = [ ( -186 ) + 2 ] = [ -184 ] = -184

Directrix (directrix):

Work from the inside out: Denominator: [ 7 * ( 8 - 2 ) + 4] [ 7 * ( 8 - 2 ) + 4] = [ 7 * ( 6 ) + 4] = [ 42 + 4] = [ 46 ] = 46

Directrix (directrix):

[ 6 * [ 5 * ( 3 - 9 ) -1 ]+ 2 ] / [ 7 * ( 8 - 2 ) + 4] = ( -184) / 46 = ?